# Essential Dimension of Central Simple Algebras when the Characteristic is Bad

### v1

@@ 1. Introduction
Fix a prime number $p$ and an algebraically closed field $k$ of $\operatorname{char}(k)=p$. Let $\text{Fields}_{k}$ be the category of fields containing $k$. Given a covariant functor $\mathcal{F}$ from $\text{Fields}_{k}$ to the category of sets, we define the essential dimension of any $A$ in $\mathcal{F}(K)$, denoted $\operatorname{ed}(A)$, to be the minimal possible $d$ for which there exists $E\in\text{Fields}_{k}$ with $E\subseteq K$ such that $\operatorname{tr.deg}(E/k)=d$ and $A=B_{K}$ where $B\in\mathcal{F}(E)$. The essential dimension of $\mathcal{F}$, denoted $\operatorname{ed}(\mathcal{F})$, is the supremum on the essential dimension of $A$ where $A$ ranges over all elements in $\mathcal{F}(K)$ and $K$ ranges over all the fields containing $k$. The essential $p$-dimension of $A$ in $\mathcal{F}(K)$, denoted $\operatorname{ed}_{p}(A)$, is defined to be the minimum on the essential dimension of $A_{L}$ where $L$ ranges over all field extensions of $K$ of degree prime to $p$. The essential $p$-dimension of $\mathcal{F}$ is thus defined to be the supremum on the essential $p$-dimension of $A$ where $A$ ranges over all elements in $\mathcal{F}(K)$ and $K$ ranges over all the fields containing $k$. In particular, $\operatorname{ed}_{p}(\mathcal{F})\leqslant\operatorname{ed}(\mathcal{F})$.
@@
The main functor under discussion in this survey is $\operatorname{Alg}_{d,e}$ in which is the functor mapping each $K$ to the set of isomorphism classes of central simple algebras of degree $d$ over $K$ and exponent dividing $e$. One is said to be studying the ‘bad characteristic case’ if $p=\operatorname{char}(k)>0$ divides $d$, the degree of the central simple algebras.
@@
The notion of essential dimension was initially defined by Reichstein and introduced in a series of seminal papers, including [BR97], [Rei00] and [RY00] (the latter is where the essential $p$-dimension was introduced). The essential dimension of $\operatorname{Alg}_{d,e}$ has received substantial attention in the literature, mostly due to its connection to linear algebraic groups of type A. But in general, its value is known only in a handful of cases, and in most cases we are very far from knowing its precise value.
A big survey on the essential dimension of various linear algebraic groups was carried out by Merkurjev [Mer13] and again in [Mer17]. The previous year, Garibaldi and Guralnick published their own work providing upper bounds for $\operatorname{Alg}_{d,e}$ in certain cases [GG16], improving results from [Bae11], [BM12] and [CM13]. A survey dedicated to central simple algebras, including a section on essential dimension, was carried out by Auel, Brussel, Garibaldi and Vishne [ABGV11].
There are several reasons to write a new survey on this topic:
  • The previous surveys on related topics were written about a decade ago, and there is a significant body of work not covered there.
  • There has not been any survey dedicated specifically to $\operatorname{Alg}_{d,e}$ in the bad characteristic ([GG16] does not focus on this functor specifically).
  • This survey serves also as an opportunity to publish a few minor results and improvements of existing results that require such a frame to be included in a publication.
@@
The survey is organized as follows: we begin with the basic background on central simple algebras and Brauer groups. Then, we introduce the description of the Brauer group in terms of differential forms (which produce the Kato-Milne cohomology groups analogous to the Galois cohomology groups in the usual case). We deduce lower bounds for the essential dimension from this setting. We then turn to the known upper bounds for the essential dimension of $\operatorname{Alg}_{2^{n},2}$, with a new upper bound for $\operatorname{Alg}_{8,2}$. A part of this survey (Section 6) is dedicated to the essential dimension of sequences of linked cyclic $p$-algebras, where we cover the known results from the literature, in addition to translations of existing results not originally set in the context of essential dimension, to the language of essential dimension (see Section 10 for the application of an observation from [CQM24] to the essential dimension of tensor products of cyclic algebras of degree $p$). We also include a few words on the essential dimension of the functor $\operatorname{Dec}_{p^{m},n}$ and extend certain results from [McK17] to the non-prime case.
@@
Table 1 records the known essential dimension bounds and values in the bad characteristic case. In the first column the functors are defined over fields containing a fixed field $k$ of characteristic $p$. When present, s|rs\,|\,r and $p|s$ while $m$ and $n$ are positive integers. Relevant citations are given in the third column and the last column records the section of this survey where the result is discussed or improved.
@@
Table 1. Essential Dimension values and bounds in the bad characteristic case
Functor Bound or Value Relevant Citations Section
$\operatorname{Alg}_{n,n}$ $\leqslant n^{2}-3n+1$ [GG16] 4
$\operatorname{ed}(\operatorname{Alg}_{p^{n},p})$ $\geqslant n+1$ [McK17], [CM20] 8, 10
$\operatorname{ed}(H^{1}_{p})$, $\operatorname{ed}(H^{1}_{p^{2}})$ 1, 2 (resp) [Led04] 8
$\operatorname{ed}_{p}(H^{1}_{p^{n}})$ 1 [RV18], this survey when $n=2$ 8
$\operatorname{ed}(\operatorname{Alg}_{r,s})$ $\geqslant\operatorname{ed}(\operatorname{SL}_{r}/\mu_{s})-1$ [Bae12, Remark 1 following 6.1] 11
$\operatorname{ed}(\operatorname{Alg}_{2n,2})$, $n>4$ even $\leqslant 2n^{2}-3n-6$ [GG16, Thm 1.3] 11, 10
$\operatorname{ed}_{p}(\operatorname{Alg}_{p,p}))$ 2 [RY00, Lem 8.5 (7)] 9
$\operatorname{ed}((\mathbb{Z}/p\mathbb{Z})^{\times n})$ 1 [Led04] 10
$\operatorname{ed}_{2}(\operatorname{Alg}_{2^{n},2})$ $\leqslant 2^{2n-2}$ [Bae12] 10
$\operatorname{ed}(\operatorname{Alg}_{4,2})$ 3 [McK17] & this survey 4
$\operatorname{ed}(\operatorname{Alg}_{4,4})$ $\geqslant 4$ and $\leqslant 5$ [Bae17] 4
$\operatorname{ed}(\operatorname{Alg}_{8,2})$ $\geqslant 4$ and $\leqslant 8$ [McK17] & this survey (resp) 5
$\operatorname{ed}(\operatorname{LCA}_{p,2})$ 3 [Cha24] 6
$\operatorname{ed}(\operatorname{LCA}_{2,n})$, $\operatorname{ed}_{2}(\operatorname{LCA}_{2,n})$ $n+1$ this survey 6
$\operatorname{ed}_{p}(\operatorname{Dec}_{p^{m},n})$ $n+1$ this survey 7
@@ 2. Central Simple Algebras
Wedderburn proved that each central simple algebra $A$ over $F$ decomposes as $D\otimes M_{t}(F)$ where $D$ is a division algebra with center $F$. This $D$ is also a central simple $F$-algebra, and thus has a degree. The index of $A$, denoted $\operatorname{ind}(A)$, is defined to be the degree of $D$.
@@
One important aspect of central simple algebras is that if the degree $d$ of $A$ is $m\cdot n$ with $\gcd(m,n)=1$, then $A\cong B\otimes C$ where $B$ and $C$ are central simple $F$-algebras with $\deg B=m$ and $\deg C=n$. This motivates the study of the different $p^{m}$-torsions of ${\operatorname{Br}}(F)$ for different prime integers $p$ and positive integers $m$. These subgroups are denoted $_{p^{m}}{\operatorname{Br}}(F)$. The case where $p$ is also the characteristic of the field $F$ is unique in many respects, and this survey is dedicated to this specific case. Algebras whose Brauer class is in $_{p^{m}}{\operatorname{Br}}(F)$ with $\operatorname{char}(F)=p$ are called “$p$-algebras”. The structure of $p$-algebras was developed by Albert in the 1930’s, culminating with the result that every $p$-algebra is ‘cyclically representable’, that is, equivalent in ${\operatorname{Br}}(F)$ to a cyclic algebra ([Alb68, Chapter VII, Theorem 31]).
There is a vast literature on central simple algebras, but the best references for the general theory are Gille and Szamuely’s book [GS17], Berhuy and Oggier’s book [BO13], Rowen’s book [Row08, Ch. 24] and to some extent Albert’s book [Alb68], although the last is an old reference and much of the terminology has changed since its publication.
3. Finitely Generated Fields
We take here the opportunity to say a few words on fields of finite transcendence degree $n$ over an algebraically closed field $k$ with $\operatorname{char}(k)=p>0$. Such fields are known to be $C_{n}$-fields, which means that every homogeneous polynomial equation of degree $d$ in more than $d^{n}$ variables has a nontrivial solution (see [EKM08, Section 97]). Since $\operatorname{char}(k)=p$, the cohomological $p$-dimension of these fields is $\leqslant n$ ([AB10]), which means $H_{p}^{n+1}(F)=0$ (see Section 8 below). In particular, when $p=2$, this means that all quadratic $(n+1)$-fold Pfister forms over such fields are hyperbolic. This particular phenomenon is used in Proposition 6.3 to prove that the essential dimension of linked sequences of $n$ quaternion algebras is exactly $n+1$.
Another aspect that is used in the literature and in this survey is the $p$-rank of a field. When $F$ is a finitely generated field containing $k$ (and thus $\operatorname{char}(F)=p$), $F$ is a vector space over $F^{p}$ of dimension $p^{r}$ for some $r<\infty$, and this $r$ is called the “$p$-rank” of $F$. When $F$ is of transcendence degree $n$ over $k$, its $p$-rank is $n$ (see [Bou90, Chapter V, Section 16.6, Corollary 3]). This connection plays a role in providing a lower bound for the essential dimension of $\operatorname{Alg}_{p^{n},p}$ in Section 10.
@@ 4. Algebras of Degree 4
A nonsplit algebra of degree 4 is either of exponent 4 or 2. The case of algebras of exponent 2 is well-understood. In this case, every central simple algebra $A$ of degree 4 and exponent 2 over $F$ is isomorphic to a tensor product of two quaternion algebras (see [Alb68, Page 174, Theorem 2]). On the one hand, the essential dimension of $A$ is at most 3 (see [McK17]). And on the other hand, if $A$ is a division algebra (for example, the algebra $\alpha^{-1},\beta)_{2,F}\otimes[\gamma,\alpha)_{2,F}$ over $F=k(\alpha,\beta,\gamma)$, see [Cha20]), then it cannot descend to a field of transcendence 2 over an algebraically closed field, because such fields are linked, i.e., every two quaternion algebras share a common maximal subfield, and then every biquaternion algebra over such fields are never division. Therefore $\operatorname{ed}(\operatorname{Alg}_{4,2})=3$.
In [Bae17], it was shown that $4\leqslant\operatorname{ed}(\operatorname{Alg_{4,4}})\leqslant 5$. The main part of this paper, which was to prove the lower bound, is based on the invariant from [Tig06] that assigns a quadratic 4-fold Pfister form to any central simple algebra of degree 4 over a field of characteristic 2, which is hyperbolic if and only if the underlying algebra is cyclic. If the essential dimension of $\operatorname{Alg}_{4,4}$ were $\leqslant 3$, that would produce a 4-fold Pfister form over a field of transcendence degree $\leqslant 3$ over an algebraically closed field. Such forms are always hyperbolic (because such fields are of cohomological 2-dimension $\leqslant 3$), and that would mean that every degree 4 algebra is cyclic. However, non-cyclic algebras of degree 4 exist (see [Alb32]). Therefore, $\operatorname{ed}(\operatorname{Alg}_{4,4})\geqslant 4$. The upper bound for the essential dimension of $\operatorname{Alg}_{n,n}$ from [GG16, Theorem 1.4] is $n^{2}-3n+1$, which for $n=4$ becomes 5.
@@ 5. Algebras of Degree 8 and Exponent 2
The essential dimension of $\operatorname{Alg}_{8,2}$ has yet to be precisely evaluated. In [Bae11], a lower bound of $3$ was provided, which was improved to 4 in [McK17]. An upper bound of 10 was provided in [Bae11], but the argument contained an error (which, if fixed, increases the upper bound to 12). This bound was improved to 9 in [Cha23]. The computation made use of the essential dimension of $(\mathbb{Z}/p\mathbb{Z})^{\times n}$ over fields of characteristic $p$, which again is 1 ([Led04]).
In this section, we recall the discussion from [Cha23] on the essential dimension of central simple algebras of degree $8$ and exponent 2 over fields of characteristic 2, and improve the obtained upper bound with a small modification of the argument, which avoids using the essential dimension of $(\mathbb{Z}/p\mathbb{Z})^{\times n}$.
Lemma 5.1 (cf. [Bae11, Lemma 3.3]).
@@
If $Q$ is split, the statement follows trivially. Suppose $Q$ is not split. The corestriction of $Q$ is trivial, which means $a,b^{2}\alpha+bc+c^{2})_{2,F}$ is split (see [MM91] for background on the corestriction of symbol algebras in the bad characteristic). Therefore, $a\equiv x^{2}(b^{2}\alpha+bc+c^{2})\pmod{\wp(F)}$ for some $x\in F^{\times}$. Hence, $Q=[a,bi+c)_{2,K}=[x^{2}(b^{2}\alpha+bc+c^{2}),bi+c)_{2,K}=[x^{2}(b^{2}(i^{2}+i)+bc+c^{2}),bi+c)_{2,K}=[xbi+x^{2}b^{2}i+x^{2}bc+xc),bi+c)_{2,K}=[(bi+c)(x+x^{2}b),bi+c)_{2,K}=[(bi+c)(x+x^{2}b),x+x^{2}b)_{2,K}=[a,x+x^{2}b)_{2,K}$. ∎
@@
The essential dimension of central simple algebras of degree 8 and exponent 2 over fields of characteristic 2 is at most 8.
@@
Given such an algebra $A$ over a field $F$ containing the small field $k$ in the background, we know that $A$ contains a subfield $F[x,y,z:x^{2}+x=\alpha,y^{2}+y=\beta,z^{2}+z=\gamma]$ for some $\alpha,\beta,\gamma\in F$ by [Row84]. The centralizer $B$ of $L=F[x]$ in $A$ is thus $\beta,ax+b)_{2,L}\otimes[\gamma,cx+d)_{2,L}$. If $a=0$, the algebra $A$ decomposes as a tensor product of three quaternion algebras, and this case was covered already in [McK17]. Suppose $a\neq 0$. Then, by replacing $x$ with $x+\frac{b}{a}$, one may assume that $b=0$. Now, the corestriction of $B$ back to $F$ is trivial, so $\beta,\alpha a^{2})_{2,F}=[\gamma,\alpha c^{2}+cd+d^{2})_{2,F}$. By the chain lemma for quaternion algebras, there exists $e\in F$ for which $\beta,\alpha a^{2})_{2,F}=[e,\alpha a^{2})_{2,F}=[e,\alpha c^{2}+cd+d^{2})_{2,F}=[\gamma,\alpha c^{2}+cd+d^{2})_{2,F}$. Therefore, there exist $r,s,t,\lambda,\tau,\theta\in F$ for which $\beta+e=\lambda^{2}+\lambda+r^{2}a^{2}\alpha$, $e=\tau^{2}+\tau+s^{2}\alpha a^{2}(\alpha c^{2}+cd+d^{2})$ and $\gamma+e=\theta^{2}+\theta+t^{2}(\alpha c^{2}+cd+d^{2})$. By replacing $e$ with $e+\tau^{2}+\tau$, $\beta$ with $\beta+\lambda^{2}+\lambda+\tau^{2}+\tau$ and $\gamma$ with $\gamma+\theta^{2}+\theta+\tau^{2}+\tau$, we get rid of the occurrences of $\lambda,\tau,\theta$. Suppose $r,s,t\neq 0$ (when one of the elements $r,s,t$ is zero, the argument is easier, see [Cha23]). So
@@
and $A\sim_{Br}[\alpha,f)_{2,F}\otimes[\beta+e,\frac{1}{r}+a)_{2,F}\otimes[\gamma+e,\frac{1}{t}+c)_{2,F}\otimes[e,\frac{1}{s}+ac+ad)_{2,F}$ (the descent computations here follow Lemma 5.1). The proof of why the last tensor product of 4 quaternion algebras is actually of index 8 is the same as in [Cha23], and thus $A$ descends to a central simple algebra of degree 8 and exponent 2 over the field $k(\alpha,\beta,\gamma,a,c,d,r,f)[s,t:s^{2}\alpha a^{2}(\alpha c^{2}+cd+d^{2})=\beta+r^{2}a^{2}\alpha=\gamma+t^{2}(\alpha c^{2}+cd+d^{2})]$, which is of transcendence degree at most 8 over $k$. ∎
As pointed out to us by J.P. Tignol, despite the two existing papers on this question ([Bae11] and [Cha23]), if one were to follow the argument of [Row84, Theorem 2’] carefully, one would probably obtain an argument very similar to our Theorem 5.2, with the same upper bound.

### v2

@@ 1. Introduction
Fix a prime number $p$ and an algebraically closed field $k$. Let $\text{Fields}_{k}$ be the category of fields containing $k$. Given a covariant functor $\mathcal{F}$ from $\text{Fields}_{k}$ to the category of sets, we define the essential dimension of any $A$ in $\mathcal{F}(F)$, denoted $\operatorname{ed}(A)$, to be the minimal possible $d$ for which there exists $E\in\text{Fields}_{k}$ with $E\subseteq F$ such that $\operatorname{tr.deg}(E/k)=d$ and $A=B_{F}$ where $B\in\mathcal{F}(E)$. The essential dimension of $\mathcal{F}$, denoted $\operatorname{ed}(\mathcal{F})$, is the supremum on the essential dimension of $A$, where $A$ ranges over all elements in $\mathcal{F}(F)$ and $F$ ranges over all the fields containing $k$. The essential $p$-dimension of $A$ in $\mathcal{F}(F)$, denoted $\operatorname{ed}_{p}(A)$, is defined to be the minimum on the essential dimension of $A_{L}$, where $L$ ranges over all field extensions of $F$ of degree prime to $p$. The essential $p$-dimension of $\mathcal{F}$ is thus defined to be the supremum on the essential $p$-dimension of $A$, where $A$ ranges over all elements in $\mathcal{F}(F)$ and $F$ ranges over all the fields containing $k$. In particular, $\operatorname{ed}_{p}(\mathcal{F})\leqslant\operatorname{ed}(\mathcal{F})$.
The essential $p$-dimension is closely linked to the concept of $p$-special fields. Recall that a field $F$ is $p$-special if it has no finite extension fields of degree prime to $p$ [22, Section 101.B]. In [41] a somewhat different definition of a $p$-special field appears: a field which has only field extensions of degree a power of $p$. The definitions coincide (see [33, Lemma 2.1]). Thus, a $p$-special field $K$ is obtained as the minimal field extension of $F$ inside its algebraic closure $\bar{F}$ that has no proper prime-to-$p$ extension. In particular, $\operatorname{ed}_{p}(A)=\operatorname{ed}(A_{K})$ for any $A\in\mathcal{F}(F)$.
@@
The main functor under discussion in this survey is $\operatorname{Alg}_{d,e}$ which is the functor mapping each field $F\supset k$ to the set of isomorphism classes of central simple algebras of degree $d$ over $F$ and exponent dividing $e$. One is said to be studying the ‘bad characteristic case’ if $p=\operatorname{char}(k)>0$ divides $d$, the degree of the central simple algebras and, conversely, we are in the ‘good characteristic’ case if $p\nmid d$.
@@
The notion of essential dimension was initially defined by Reichstein and introduced in a series of seminal papers, including [12], [42] and [44] (the latter is where the essential $p$-dimension was introduced). The essential dimension of $\operatorname{Alg}_{d,e}$ has received substantial attention in the literature, mostly due to its connection to linear algebraic groups of type A. But in general, its value is known only in a handful of cases, and in most cases we are very far from knowing its precise value.
There have been several surveys written on essential dimension over the last 15 years. Comprehensive surveys on the essential dimension of various linear algebraic groups were written by Merkurjev [37], [38] and again in [39]. A survey on central simple algebras, including a section on essential dimension, was written by Auel, Brussel, Garibaldi and Vishne [4].
This survey focuses on the bounds and values of the essential dimension of $\operatorname{Alg}_{d,e}$ and some related functors in the bad characteristic. The essential dimension in the bad case is often different from the good case. For example, if one considers the functor $\mathcal{F}={{}_{p}{\operatorname{Br}}}$ mapping a field $F$ to the $p$-torsion $_{p}{\operatorname{Br}}(F)$} of $F$, then its essential $p$-dimension is infinite in the good case, but in the bad case, it is equal to 2. This is a result of the fact that in the bad case, every algebra of exponent $p$ is Brauer equivalent to a cyclic algebra of degree $p^{m}$ for some $m$, and that the essential $p$-dimension of $H_{p^{m}}^{1}$ in the bad case is 1. See Section 7 for a description of the functor $H^{1}_{p^{m}}$ and [40, Props 7.2, 9.2] for further details of the argument.
@@
The survey is organized as follows: we begin with the basic background on central simple algebras and Brauer groups. Then, we introduce the description of the Brauer group in terms of differential forms (which produce the Kato-Milne cohomology groups analogous to the Galois cohomology groups in the usual case). We deduce lower bounds for the essential dimension from this setting. We then turn to the known upper bounds for the essential dimension of $\operatorname{Alg}_{2^{n},2}$, with a new upper bound for $\operatorname{Alg}_{8,2}$.
Part of this survey (Section 5) is dedicated to the essential dimension of sequences of linked cyclic $p$-algebras, where we cover the known results from the literature, in addition to translations of existing results not originally set in the context of essential dimension, to the language of essential dimension (see Section 9 for the application of an observation from [21] to the essential dimension of tensor products of cyclic algebras of degree $p$). We also include a few words on the essential dimension of the functor $\operatorname{Dec}_{p^{m},n}$ and extend certain results from [36] to the non-prime case.
@@
Table 1 records the known essential dimension bounds and values in the bad characteristic case with the corresponding good characteristic bounds provided as a comparison. In the first column the functors are defined over fields containing a fixed field $k$ of characteristic $p$ and $m$ and $n$ are positive integers. Relevant citations are given in the third column and the last column indicates the section(s) of this survey where the result is discussed or improved.
@@
Table 1. Essential Dimension values and bounds in the bad and good characteristic cases.
Functor bound/value
in bad char
bound/value
in good char
Citations Sec.
$\operatorname{ed}(\operatorname{Alg}_{n,n})$ $\leqslant n^{2}-3n+1$ same [23, 1.4],
[32]
3
$\operatorname{ed}(\operatorname{Alg}_{4,2})$,
$\operatorname{ed}_{2}(\operatorname{Alg}_{4,2})$
3 4 [5, Thm 1.1],
[8, Rmk 8.2]
3
$\operatorname{ed}(\operatorname{Alg}_{4,4})$,
$\operatorname{ed}_{2}(\operatorname{Alg}_{4,4})$
$\geqslant 4$ and $\leqslant 5$ 5 [7],
[38, Page 470]
3
$\operatorname{ed}(\operatorname{Alg}_{8,2})$,
$\operatorname{ed}_{2}(\operatorname{Alg}_{8,2})$
$\geqslant 4$ and $\leqslant 8$ 8 [36] &
this survey (resp),
[8, Cor 8.3]
4
$\operatorname{ed}((\mathbb{Z}/p\mathbb{Z})^{\times n})$ 1 $n$ [31] 4, 9
$\operatorname{ed}(\operatorname{LCA}_{p,2})$ 3 open [16] 5
$\operatorname{ed}(\operatorname{LCA}_{2,n})$,
$\operatorname{ed}_{2}(\operatorname{LCA}_{2,n})$
$n+1$ same this survey 5
$\operatorname{ed}_{p}(\operatorname{Dec}_{p^{m},n})$ $n+1$ $2n$ this survey,
[38, Ex 3.7]
6
$\operatorname{ed}_{p}(\operatorname{Alg}_{p^{n},p})$ $\geqslant n+1$ $\geqslant 2n$ [36],
[20],
[38, Ex 3.7]
7, 9
$\operatorname{ed}(H^{1}_{p})$,
$\operatorname{ed}(H^{1}_{p^{2}})$
1, 2 (resp) 1 [31] 7
$\operatorname{ed}_{p}(H^{1}_{p^{n}})$ 1 1 [43],
this survey
when $n=2$
7
$\operatorname{ed}_{p}(\operatorname{Alg}_{p,p})$ 2 same [44, Lem 8.5 (7)] 8
$\operatorname{ed}(\operatorname{Alg}_{2n,2})$,
$n>4$ even
$\leqslant 2n^{2}-3n-6$ $\leqslant 2n^{2}-3n-4$ [23, Thm 1.3] 9
$\operatorname{ed}_{2}(\operatorname{Alg}_{2^{n},2})$ $\leqslant 2^{2n-2}$ $2^{2n-4}+2^{n-1}$ [6, Thm 1.1] 9
@@ 2. Central Simple Algebras
Wedderburn proved that given a central simple algebra $A$ over $F$ there exists a division algebra $D$ with center $F$, unique up to isomorphism, and positive integer $t$ so that $A$ decomposes as $A\cong D\otimes M_{t}(F)$. $D$ is also a central simple $F$-algebra, and thus has a degree. The index of $A$, denoted $\operatorname{ind}(A)$, is defined to be the degree of $D$.
@@
One important aspect of central simple algebras is that if the degree $d$ of $A$ is $m\cdot n$ with $\gcd(m,n)=1$, then $A\cong B\otimes C$ where $B$ and $C$ are central simple $F$-algebras with $\deg B=m$ and $\deg C=n$. This motivates the study of the $p^{m}$-torsion subgroups of ${\operatorname{Br}}(F)$ for different prime integers $p$ and positive integers $m$. These subgroups are denoted $_{p^{m}}{\operatorname{Br}}(F)$. Though this notion is independent of the characteristic, we focus on the case where $p$ is also the characteristic of the field $F$ as this survey is dedicated to the so-called bad characteristic case, which is unique in many respects. Algebras whose Brauer class is in $_{p^{m}}{\operatorname{Br}}(F)$ with $\operatorname{char}(F)=p$ are called “$p$-algebras”. The structure of $p$-algebras was developed by Albert in the 1930s, culminating with the result that every $p$-algebra is ‘cyclically representable’, that is, equivalent in ${\operatorname{Br}}(F)$ to a cyclic algebra ([1, Chapter VII, Theorem 31]).
There is a vast literature on central simple algebras, several references for the general theory include Gille and Szamuely’s book [25], Berhuy and Oggier’s book [10], Rowen’s book [46, Ch. 24] and to some extent Albert’s book [1], although the last is an old reference and much of the terminology has changed since its publication.
@@ 3. Algebras of Degree 4
In this section, $p=2$, and so $k$ is an algebraically closed field of $\operatorname{char}(k)=2$. For a field $F$ containing $k$, any central simple $F$-algebra of degree 4 is either of exponent 4, 2 or 1 (the split case). The case of algebras of exponent 2 is well-understood.
Proposition 3.1.
The essential dimension and essential 2-dimension of $\operatorname{Alg}_{4,2}$ is exactly 3.
Proof.
Let $A$ be a central simple algebra of degree 4 and exponent 2 over $F$. By [1, Page 174, Theorem 2], $A$ is isomorphic to a tensor product of two quaternion algebras. On the one hand, the essential dimension of $A$ is at most 3 by [36]. On the other hand, let $A$ be a division algebra over $F$. For example, the algebra $\alpha^{-1},\beta)_{2,F}\otimes[\gamma,\alpha)_{2,F}$ over $F=k(\alpha,\beta,\gamma)$ is division (see [14]). $A$ remains division after any odd degree extension and any such extension cannot descend to a field of transcendence degree 2 over an algebraically closed field, because such fields are linked (see [9, Theorem 3.1] where the statement is about fields of $u$-invariant $\leqslant 4$ which applies because the $u$-invariant in this case is 4 by [35, Prop 2]), i.e., every two quaternion algebras share a common maximal subfield, and then every biquaternion algebra over such fields is never a division algebra. Therefore $\operatorname{ed}(\operatorname{Alg}_{4,2})=3$.
Compare this result to the lower bound on the essential dimension of $\operatorname{Alg}_{p^{n},p}$ given in Proposition 9.2. In [7], it was shown that $4\leqslant\operatorname{ed}(\operatorname{Alg_{4,4}})\leqslant 5$. Here is a short explanation of this result:
Proposition 3.2.
The essential dimension and essential 2-dimension of $\operatorname{Alg}_{4,4}$ are at least 4 and at most 5.
Proof.
There is the invariant from [47] that assigns a quadratic 4-fold Pfister form to any central simple algebra of degree 4 over a field of characteristic 2, which is hyperbolic if and only if the underlying algebra is cyclic. If the essential dimension of $\operatorname{Alg}_{4,4}$ were $\leqslant 3$, that would produce a 4-fold Pfister form over a field of transcendence degree $\leqslant 3$ over an algebraically closed field. Such forms are always hyperbolic (because the $u$-invariant of such fields is at most $8$), and that would mean that every degree 4 algebra is cyclic. However, non-cyclic algebras of degree 4 exist (see [2]). Therefore, $\operatorname{ed}(\operatorname{Alg}_{4,4})\geqslant 4$. For $\operatorname{ed}_{2}(\operatorname{Alg}_{4,4})\geqslant 4$ it is enough to add that anisotropic quadratic 4-fold Pfister forms remain anisotropic under odd extensions.
The upper bound for the essential dimension of $\operatorname{Alg}_{n,n}$ from [23, Theorem 1.4] is $n^{2}-3n+1$, which for $n=4$ becomes 5. Since $\operatorname{ed}_{2}(\operatorname{Alg}_{4,4})\leqslant\operatorname{ed}(\operatorname{Alg}_{4,4})$, the upper bound for the essential 2-dimension follows as well. ∎
@@ 4. Algebras of Degree 8 and Exponent 2
The essential dimension of $\operatorname{Alg}_{8,2}$ has yet to be precisely evaluated. In [5], a lower bound of $3$ was provided, which was improved to 4 in [36]. An upper bound of 10 was provided in [5], but the argument contained an error (which, if fixed, increases the upper bound to 12). This bound was improved to 9 in [15]. The computation made use of the essential dimension of $(\mathbb{Z}/p\mathbb{Z})^{\times n}$ over fields of characteristic $p$, which is equal to 1 ([31]). The definition of the essential dimension of a finite group appears in [12] and one can consult [31] for the specific case of a finite $p$-group over a field of characteristic $p$. In short, the essential dimension of a finite group $G$ is the essential dimension of Galois extensions of $F$ with Galois group $G$, and in particular, $\operatorname{ed}(\mathbb{Z}/p^{m}\mathbb{Z})=\operatorname{ed}(H^{1}_{p^{m}})$ (see Section 7).
In this section, we recall the discussion from [15] on the essential dimension of central simple algebras of degree $8$ and exponent 2 over fields of characteristic 2, and improve the upper bound from [15] with a small modification of the argument, which avoids using the essential dimension of $(\mathbb{Z}/p\mathbb{Z})^{\times n}$.
Lemma 4.1 (cf. [5, Lemma 3.3]).
@@
If $Q$ is split, the statement follows trivially. Suppose $Q$ is not split. The corestriction of $Q$ is trivial, which means $a,b^{2}\alpha+bc+c^{2})_{2,F}$ is split (see [34] for background on the corestriction of symbol algebras in the bad characteristic). Therefore, $a\equiv x^{2}(b^{2}\alpha+bc+c^{2})\pmod{\wp(F)}$ for some $x\in F^{\times}$. Hence, $Q=[a,bi+c)_{2,K}=[x^{2}(b^{2}\alpha+bc+c^{2}),bi+c)_{2,K}=[x^{2}(b^{2}(i^{2}+i)+bc+c^{2}),bi+c)_{2,K}=[xbi+x^{2}b^{2}i+x^{2}bc+xc),bi+c)_{2,K}=[(bi+c)(x+x^{2}b),bi+c)_{2,K}=[(bi+c)(x+x^{2}b),x+x^{2}b)_{2,K}=[a,x+x^{2}b)_{2,K}$. ∎
@@
The essential dimension of a central simple algebra of degree 8 and exponent 2 over a field of characteristic 2 is at most 8.
@@
Given such an algebra $A$ over a field $F$ containing the base field $k$ of characteristic 2, we know that $A$ contains a subfield $F[x,y,z:x^{2}+x=\alpha,y^{2}+y=\beta,z^{2}+z=\gamma]$ for some $\alpha,\beta,\gamma\in F$ by [45]. The centralizer $B$ of $L=F[x]$ in $A$ is thus $\beta,ax+b)_{2,L}\otimes[\gamma,cx+d)_{2,L}$. If $a=0$, the algebra $A$ decomposes as a tensor product of three quaternion algebras, and this case was covered already in [36]. Suppose $a\neq 0$. Then, by replacing $x$ with $x+\frac{b}{a}$, one may assume that $b=0$. Now, the corestriction of $B$ back to $F$ is trivial, so $\beta,\alpha a^{2})_{2,F}=[\gamma,\alpha c^{2}+cd+d^{2})_{2,F}$. By the chain lemma for quaternion algebras, there exists $e\in F$ for which $\beta,\alpha a^{2})_{2,F}=[e,\alpha a^{2})_{2,F}=[e,\alpha c^{2}+cd+d^{2})_{2,F}=[\gamma,\alpha c^{2}+cd+d^{2})_{2,F}$. Therefore, there exist $r,s,t,\lambda,\tau,\theta\in F$ for which $\beta+e=\lambda^{2}+\lambda+r^{2}a^{2}\alpha$, $e=\tau^{2}+\tau+s^{2}\alpha a^{2}(\alpha c^{2}+cd+d^{2})$ and $\gamma+e=\theta^{2}+\theta+t^{2}(\alpha c^{2}+cd+d^{2})$. By replacing $e$ with $e+\tau^{2}+\tau$, $\beta$ with $\beta+\lambda^{2}+\lambda+\tau^{2}+\tau$ and $\gamma$ with $\gamma+\theta^{2}+\theta+\tau^{2}+\tau$, we get rid of the occurrences of $\lambda,\tau,\theta$. Suppose $r,s,t\neq 0$ (when one of the elements $r,s,t$ is zero, the argument is easier, see [15]). So
@@
and $A\sim_{Br}[\alpha,f)_{2,F}\otimes[\beta+e,\frac{1}{r}+a)_{2,F}\otimes[\gamma+e,\frac{1}{t}+c)_{2,F}\otimes[e,\frac{1}{s}+ac+ad)_{2,F}$ (the descent computations here follow Lemma 4.1). The proof of why the last tensor product of 4 quaternion algebras is actually of index 8 is the same as in [15], and thus $A$ descends to a central simple algebra of degree 8 and exponent 2 over the field $k(\alpha,\beta,\gamma,a,c,d,r,f)[s,t:s^{2}\alpha a^{2}(\alpha c^{2}+cd+d^{2})=\beta+r^{2}a^{2}\alpha=\gamma+t^{2}(\alpha c^{2}+cd+d^{2})]$, which is of transcendence degree at most 8 over $k$. ∎
It was pointed out to us by J.P. Tignol that the argument presented in the proof of 4.2 follows closely to the argument of Rowen’s in [45, Theorem 2’].